OpenStudy (anonymous):

The area of a rectangular swimming pool is given by 3x^2-2x-1ft^2 . One side length of the ool is given by 3x + 1 feet. What is an lgebrapic expression for the other side length of the apool? Simplify this, and include correct units as part of your answer.

OpenStudy (anonymous):

how do I find something like this for an answer

OpenStudy (anonymous):

Area= L * W Area ------ = W L you know area and lenght

OpenStudy (anonymous):

3x^2-2x-1 ----------- = other sides 3x+1

OpenStudy (anonymous):

factored-----> (3x + 1)(x - 1) ----------------- (3x+1) (x-1) is other sides

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Thanks that was easy to follow and merry christmas

OpenStudy (anonymous):

you too